. A furnace uses a coal the ultimate analysis shows 57.22 % C, 5.64 % H, 1.41 % N, 5.06 % S, and 14.67 % ash. If the sulfur in the ash is negligible and the excess air is 80 %, compute the Orsat analysis of the combustion gases.









2.) If dry hydrogen gas at atmospheric pressure is burnt completely in a furnace with 32 % excess air, calculate the Orsat analysis of the flue gas from this furnace.




1
Expert's answer
2022-01-21T09:48:01-0500

1. 2H2 + O2 = 2H2O

100 x 0.96 = 96 kmols H2.

96/20 x 0.68= 70,5882353 kmols O2.

4 kmols H2O.

2. 2CO + O2 = 2CO2

2H2 + O2 = 2H2O

CH4 + 2O2 = CO2 + 2H2O

C2H6 + 2.5O2 = CO2 + 3H2O

50/2= 25 O2. 50 CO2.

30 /2= 15 O2. 30 H2O.

5x2 = 10 O2. 5 CO2 and 10 H2O.

5x2 = 10 O2. 5 CO2 and 10 H2O.

25+15+10+10=60 kmols O2

60/0.5 = 120. 120 + 6 = 126 O2.